### v1

@@ Discussion
This paper used bottom-up modeling to estimate the physical and economic impacts of electrification of all residential appliances and personal vehicles on distribution grids in each county of the lower 48 United States. The modeling tools are open-source; the cleaned input data and simulation results are free and public. This paper found that distribution grid reinforcement requirements depend mainly on space heating demand, which can triple peak aggregate demand in the coldest counties relative to business as usual. Without demand-side management, nationwide grid reinforcements could cost \$183 to \$415 billion, or \$1,500 to \$3,400 per household, but smart electrification could reduce these costs by up to 92%.
This paper’s results could find use by grid planners, for whom spatially-resolved estimates of grid reinforcement requirements can inform capacity expansion decisions. The demand-side management results could also find use by policymakers who determine incentives and regulations that shape individuals’ electrification choices. For example, the effectiveness of device coordination in reducing grid reinforcement costs could motivate policymakers to encourage or require electrification with appliances and vehicles that have the sensing, communication, and computing capabilities that device coordination systems require.
@@
There are many opportunities to extend this work. This paper’s scope covers distribution grid reinforcement driven by electrification of the residential sector and personal vehicles. Future research could extend this work to include transmission grid reinforcement, generation capacity expansion, and electrification of the commercial and industrial sectors. In particular, estimating transmission and generation costs from residential electrification would be straightforward and useful extensions. Future research could also extend the open-source EDGIE toolbox to incorporate solar photovoltaics, stationary batteries, or thermal energy storage. Although we could not find reliable, spatially-resolved data on distribution grid topologies or reinforcement prices, such data would enable analysis of power quality issues and more precise estimation of distribution grid reinforcement costs. This paper used historical weather data; future work could incorporate the effects of climate change on demand peaks driven by extreme weather. Finally, while this paper estimated the value of demand-side management strategies in mitigating distribution grid reinforcement costs, these strategies could also reduce generation and transmission expansion costs, greenhouse gas emissions, air pollution, and household energy bills. Identifying appropriate, location-specific mixes of grid upgrades and demand-side management measures is an interesting direction for future work.
Methods
[…]
@@ Methods
We tune model parameters to match real data wherever possible. Distributions of building and appliance properties follow the ResStock database [42]. Heat pump coefficient of performance curves come from manufacturer data aggregated by the Northeast Energy Efficiency Partnerships [43]. We generate domestic hot water time series from [32]. The number of vehicles per household in each county comes from census data [44]. We classify personal vehicles as either small (e.g., coupes and sedans) or large (e.g., pickup trucks and sport utility vehicles) and tune the distribution of small and large vehicles in each county to survey data from [45]. We tune the driving efficiencies of small and large electric vehicles to the Tesla Model S and Ford F-150 lightning, respectively. For both vehicle types, driving efficiencies vary with outdoor temperatures according to [46]; this variation reflects energy use for cabin heating and cooling, as well as temperature-dependent battery chemistry. Commute distances follow census data for each county [47]. The Supplementary Information document contains further details on modeling, simulation, and data sources. Table 1 summarizes the data sources drawn upon by this paper.
@@
Parameter Data Source References
Building types and floor areas National Renewable Energy Laboratory’s ResStock [23, 48]
Business-as-usual appliances National Renewable Energy Laboratory’s ResStock [23, 48]
Building insulation and air sealing ASHRAE code [49]
Design heating and cooling temperatures EnergyStar [33]
Heat pump coefficients of performance Northeast Energy Efficiency Partnerships [43]
Domestic hot water profiles Building America analysis spreadsheets [32]
Commute times Census survey [50]
Vehicles per household Census survey [50]
Electric vehicle parameters Manufacturer specification and journals [46, 51]
Weather data Oikolab [29]
Climate zones ASHRAE/IECC [52]
Grid upgrade prices Journals [37]
[…]
Space heating and cooling EDGIE models indoor temperature dynamics using first-order linear ordinary differential equations:
@@
$p_{1}(k)=\begin{cases}0&\hat{q}_{1}(k)\leq 0\\ \hat{q}_{1}(k)/\eta_{1}(k)&0<\hat{q}_{1}(k)\leq\eta_{1}(k)\overline{p}_{1h}\\ (1-\eta_{1}(k))\overline{p}_{1h}+\hat{q}_{1}(k)&\eta_{1}(k)\overline{p}_{1h}<\hat{q}_{1}(k)\leq\overline{p}_{1h}+\overline{p}_{1r}\\ \overline{p}_{1h}+\overline{p}_{1r}&\overline{p}_{1h}+\overline{p}_{1r}<\hat{q}_{1}(k).\\ \end{cases}$ (4)
Experimental validation of the space heating model can be found in [28]. Space cooling works similarly but without backup resistance heat.
[…]
Electric vehicles EDGIE models electric vehicle batteries via
@@
[…]
Water heating EDGIE models water heaters via
@@
Declarations
[…]
Data and code availability The data and code that support this study are available at https://github.com/priyada7/EDGIE.
[…]
Supplementary information The online version contains supplementary material available at LINK.

### v2

@@ DISCUSSION
This paper used bottom-up modeling to estimate the physical and economic impacts of electrification of all residential appliances and personal vehicles on distribution grids in each county of the lower 48 United States. The modeling tools are open-source; the cleaned input data and simulation results are free and public. This paper found that distribution grid reinforcement requirements depend mainly on space heating demand, which can increase peak aggregate demand by a factor of five or more in the coldest counties. Without demand-side management, distribution grid reinforcement could cost \$320 to \$720 billion nationwide, or \$2,600 to \$5,800 per household, but smart electrification could eliminate up to three-quarters of these costs.
Sensitivity to modeling choices
This section discusses the sensitivity of nationwide distribution grid reinforcement requirements to several modeling assumptions, including heat pump efficiencies, whether households lower their heating temperature setpoints overnight, heat pump sizing protocols, and the choice of weather data. The Supplementary Information document discusses the sensitivity analyses in more detail.
Heat pump efficiencies significantly influence grid reinforcement requirements. In cold weather, heat pumps become less efficient and may require electric resistance backup to meet demand. The central cases presented above assumed that heat pumps in cold regions met the United States Department of Energy’s Cold-Climate Heat Pump Challenge specifications 34. However, in a sensitivity study where heat pumps in cold regions have efficiencies that are similar to today’s equipment, nationwide distribution grid reinforcement requirements increase by about 18%.
Lowering space heating temperature setpoints overnight can increase grid reinforcement requirements. While lowering setpoints can save energy overnight, warming the thermal mass of a building back up in the morning can require high power and potentially trigger the use of resistance backup heat. In the central cases presented above, users did not alter setpoints overnight. However, in a sensitivity study where users lower setpoints overnight, with randomized but typical setpoint adjustment magnitudes and timings, nationwide distribution grid reinforcement requirements increase by about 18%.
Heat pump sizing has a comparatively small influence on grid reinforcement requirements. In the central cases presented above, we sized heat pumps to meet the larger of the design cooling load and the design heating load, capping the size at the maximum capacity currently available on the market. If heating demand in very cold weather exceeded the heat pump’s capacity, electric resistance backup heat turned on to maintain space temperatures. An alternative approach would be to size heat pumps in colder climates to meet only the design cooling load, relying much more heavily on backup heat in cold weather. We simulated this alternative approach in a sensitivity study and found that it increases nationwide grid reinforcement requirements by about 8%. In the coldest regions, however, sizing heat pumps to the design cooling loads increases grid reinforcement requirements by about 12%.
The sensitivities of grid reinforcement requirements to modeling choices on heat pump efficiency (about 18%), overnight temperature setpoint adjustments (about 18%), and heat pump sizing (about 8%) are more than additive. In a sensitivity study including all three of the alternative modeling choices discussed above, nationwide grid reinforcement requirements increase by about 50%.
The choice of weather data for the all-electric simulations has relatively little influence on grid reinforcement requirements. In the central cases presented above, we used historical weather data from 2021. An alternative approach would be to use hypothetical future weather data that incorporates the effects of climate change. Climate scientists affiliated with the Intergovernmental Panel on Climate Change have developed frameworks known as the Shared Socioeconomic Pathways (SSP) and Representative Concentration Pathways (RCP) 42. These frameworks provide both quantitative and qualitative descriptions of various societal and environmental development pathways through 2100. For example, in the middle-of-the-road scenario SSP2-RCP4.543, global mean surface temperature rises about 2 C by 2050. In a sensitivity study that used SSP2-RCP4.5 weather data for 2050, nationwide grid reinforcement requirements decrease by about 4%. While climate change increases electricity demand peaks from air conditioning in hot regions, climate change also decreases peaks from space heating in cold regions. In the sensitivity study, the latter effect outweighs the former.
Potential impact
The findings of this paper could inform decisions about energy system design and operation. Power grid operators might use the spatially-resolved estimates of electricity demand profiles under all-electric scenarios, or the open-source software that generated them, in their capacity expansion planning. Policymakers or utility program administrators might use the demand-side management results to shape electrification incentives. For example, the perhaps surprising effectiveness of device coordination in mitigating grid impacts might warrant making electrification incentives contingent on installing equipment with the sensing, communication, and computing capabilities required for device coordination. Utility regulators might refer to the physical or economic estimates this paper when evaluating utility requests for rate increases to cover grid reinforcement costs driven by electrification. Finally, the main takeaways of this paper might influence national conversations about the future of United States energy systems.
Limitations and future work
One challenge in this study was finding reliable, spatially-resolved data on distribution grid topologies and reinforcement prices. Due to a lack of topological data, we did not analyze the impacts of electrification on power quality, such as voltage support or frequency regulation. This is one possible direction for future work. Due to a lack of spatially-resolved grid reinforcement price data, we used the same price distribution in every United States county. Higher-resolution price data could enable more precise estimation of distribution grid reinforcement costs.
@@
There are many other opportunities to extend this work. This paper’s scope covers distribution grid reinforcement driven by electrification of the residential sector and personal vehicles in the United States. Future research could extend this work to include transmission grid reinforcement and generation capacity expansion; to include electrification of the commercial and industrial sectors; or to include other countries. In particular, estimating transmission and generation costs from residential electrification would be straightforward and useful extensions. Future research could also extend the open-source EDGIE toolbox to incorporate solar photovoltaics, stationary batteries, or thermal energy storage. Finally, while this paper estimated the value of demand-side management strategies in mitigating distribution grid reinforcement costs, these strategies could also reduce transmission and generation capacity expansion costs, greenhouse gas emissions, air pollution, and household energy bills. An interesting direction for future work is to identify location-specific mixes of grid upgrades and demand-side management measures that appropriately balance the various costs and benefits.
EXPERIMENTAL PROCEDURES
@@
Table 1 summarizes the data sources drawn used in this paper. Distributions of building and appliance properties follow the ResStock database 44. Heat pump coefficient of performance curves come from manufacturer data aggregated by the Northeast Energy Efficiency Partnerships 45. We generate domestic hot water time series from 31. The number of vehicles per household in each county comes from census data 46. We classify personal vehicles as either small (e.g., coupes and sedans) or large (e.g., pickup trucks and sport utility vehicles) and tune the distribution of small and large vehicles in each county to survey data from 47. We tune the driving efficiencies of small and large electric vehicles to the Tesla Model S and Ford F-150 lightning, respectively. For both vehicle types, driving efficiencies vary with outdoor temperatures according to 48; this variation reflects energy use for cabin heating and cooling, as well as temperature-dependent battery chemistry. Commute distances follow census data for each county 49. The Supplementary Information document contains further details on modeling, simulation, and data sources.
@@
Parameter Data Source Ref.
Building types and floor areas NREL’s ResStock 23, 50
Business-as-usual appliances NREL’s ResStock 23, 50
Building insulation and air sealing ASHRAE code 51
Design heating and cooling temperatures EnergyStar 33
Heat pump coefficients of performance Northeast Energy Efficiency Partnerships 45
Domestic hot water profiles Building America analysis spreadsheets 31
Commute times Census survey 52
Vehicles per household Census survey 52
Electric vehicle parameters Manufacturer specification and journals 48, 53
Weather data Oikolab 29
Climate zones ASHRAE/IECC 54
Grid upgrade prices Journals 37
Space heating and cooling
EDGIE models indoor temperature dynamics using first-order linear ordinary differential equations:
@@
$p_{1}(k)=\begin{cases}0&\hat{q}_{1}(k)\leq 0\\ \hat{q}_{1}(k)/\eta_{1}(k)&0<\hat{q}_{1}(k)\leq\eta_{1}(k)\overline{p}_{1h}\\ (1-\eta_{1}(k))\overline{p}_{1h}+\hat{q}_{1}(k)&\eta_{1}(k)\overline{p}_{1h}<\hat{q}_{1}(k)\leq\eta_{1}(k)\overline{p}_{1h}+\overline{p}_{1r}\\ \overline{p}_{1h}+\overline{p}_{1r}&\eta_{1}(k)\overline{p}_{1h}+\overline{p}_{1r}<\hat{q}_{1}(k).\\ \end{cases}$ (4)
Experimental validation of the space heating model can be found in 28. Space cooling works similarly but without backup resistance heat.
Electric vehicles
EDGIE models electric vehicle batteries via
@@
Water heating
EDGIE models water heaters via
@@
Distribution grid reinforcement costs
Total cost calculation
Figure 7: Distribution grid reinforcement cost distribution vs. discount rate, assuming an annual inflation rate of 2.5%. The shaded gray area is the 95% confidence region.
We assume that distribution grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. Distribution grid reinforcement costs include both capital costs associated with installing new infrastructure and operation and maintenance cost increases associated with running a larger grid. The net present cost of grid reinforcement is
$\underbrace{\frac{\pi_{\text{recurring}}G}{n}\sum_{k=1}^{n}\frac{k(1+i)^{k}}{(1+r)^{k}}}_{\text{Present value of recurring costs}}+\underbrace{\frac{\pi_{\text{capital}}G}{n}\sum_{k=1}^{n}\frac{(1+i)^{k}}{(1+r)^{k}}}_{\text{Present value of capital costs}}.$ (9)
Here $\pi_{\text{recurring}}$ (\$/kW) is the initial recurring price in the first year, $\pi_{\text{capital}}$ (\$/kW) is the initial capital price in the first year, $i$ is the annual inflation rate, $r$ is the discount rate, $G$ (kW) is the total growth in grid capacity, $k$ indexes years, $n$ is the total number of years over which recurring costs are projected. Fig. 7 shows the distribution grid reinforcement cost vs. the discount rate $r$. To compute the cost estimates reported in the main document, we set the discount rate equal to the annual inflation rate of 2.5%. Higher discount rates result in lower net present cost estimates.
Grid reinforcement requirement uncertainty
Figure 8: Histogram of the all-electric peak aggregate demand of 1,000 Minnesota homes over 10,000 Monte Carlo simulations. The distribution is approximately Gaussian with a standard deviation of about 0.5% of the mean.
In principle, the grid reinforcement requirement $G$ (kW) is a random variable that depends on the randomly generated parameters in the building and machine models. However, $G$ is computed in each county by summing over 1,000 homes and about 2,000 vehicles. From the Central Limit Theorem, therefore, we expect $G$ to have an approximately Gaussian distribution with small variance. Fig. 8 shows the distribution of the peak aggregate demand for a 1,000-home neighborhood over 10,000 Monte Carlo simulations. The standard deviation of the peak aggregate demand is about 0.5% of the mean. Given that nationwide simulations require significant computing time and the variations of county-level peaks from one Monte Carlo simulation to the next are small, we conducted the 1,000-home simulations once for each county and treated the resulting grid reinforcement estimate as deterministic.
Grid reinforcement cost uncertainty
We estimate the grid reinforcement cost using the prices $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$ from 37. Since 37 does not model the uncertainty associated with these prices, we assume they follow Gaussian distributions with standard deviations of 20% of the mean prices. As the total cost in Eq. (9) is a linear combination of $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$, the total cost is also Gaussian. We calculate its mean and variance analytically using standard formulas.
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Kevin J. Kircher (kircher@purdue.edu).
Materials availability
No materials were used in this study.
Data and code availability
The data and code that support this study are available at https://github.com/priyada7/EDGIE.
ACKNOWLEDGMENTS
This work was supported through the Center for High Performance Buildings at Purdue University. The authors thank James Braun, Roy Crawford, Andrew Hjortland, Scott Munns, Wayne Craft, Ronald Domitrovic, Elias Pergantis and Sophia Evers for helpful discussion.
SUPPLEMENTAL INFORMATION INDEX
Document S1. Supplemental experimental procedures, Figures S1–S13.
AUTHOR CONTRIBUTIONS
Conceptualization: P. and K.J.K; methodology: P., C.C., K.B. and K.J.K; investigation: P.; writing – original draft: P. and K.J.K; writing – review & editing, P., C.C., K.B. and K.J.K; funding acquisition: K.J.K; resources: K.J.K.; supervision: C.C., K.B. and K.J.K.
DECLARATION OF INTERESTS
The authors declare no competing interests.